# Search for Light Dark Matter in Rare Meson Decays

### v1

@@ Abstract
Current dark matter direct detection experiments have low sensitivity to sub-GeV dark matter. In this work, we demonstrate that rare $B$ and $K$ meson decays with missing energy in the final state can serve as efficient probes in this mass range. We analyze a generic $Z^{\prime}$ portal dark matter model and derive upper limits on its parameters from experimental bounds on the rare $B$ and $K$ meson decays. Our results show that such meson decay processes provide complementary constraints to current direct detection experiments for sub-GeV dark matter, particularly for interaction forms mediated by dark matter momentum dependent operators.
@@ I Introduction
Abundant astronomical and cosmological evidence indicates that about a quarter of the matter in the Universe is composed of dark matter (DM) ParticleDataGroup:2024cfk . However, the nature of DM remains unknown as the Standard Model (SM) of particle physics fails to provide a viable DM candidate. The most widely studied DM candidates are the so-called weakly interacting massive particles (WIMPs), which could naturally provide the right amount of DM relic density, a coincidence known as the WIMP miracle. Current most DM direct detection experiments target the search for WIMPs, with mass range roughly on the order of GeV to TeV. DM direct detection experiments mostly aim to observe elastic or inelastic scatters of DM particles with atomic nuclei, such as liquid xenon detectors XENONnT XENON:2023cxc , LUX-ZEPLIN LZ:2022lsv , and PandaX-4T PandaX:2022xas . While no confirmed DM signal has been observed, these experiments have placed stringent limits on the DM–nucleon scattering cross section. However, their sensitivity drops sharply for DM masses below the GeV scale, as the nuclear recoil signals from its elastic scattering with nuclei typically lie below the detection energy threshold and are overwhelmed by background noise. Searching for light DM in heavy meson decays can serve as a complement to DM direct detection, and its sensitivity in some models may be higher than that of direct detection experiments.
@@
Semileptonic flavor-changing neutral current (FCNC) decays are highly suppressed in the SM due to their loop origin and the Glashow-Iliopoulos–Maiani (GIM) mechanism, making them exceptionally sensitive to new physics (NP) effects. In particular, decays with a neutrino pair in the final state provide a distinct probe of NP compared to those with charged leptons. For semileptonic decays to charged leptons, NP effects are primarily probed through modifications of the relevant Wilson coefficients, mediated by virtual NP particles. In contrast, for decays to a neutrino pair, NP particles themselves can exist as missing energy hiding in the final state, since the neutrinos are not detected in experiments. In addition to the SM neutrinos, any electrically neutral, undetectable particle can manifest as missing energy in such processes. Notably, the Belle II experiment has recently reported the first evidence of $B^{+}\to K^{+}\nu\bar{\nu}$ with the branching ratio of $\text{Br}(B^{+}\to K^{+}\nu\bar{\nu})=(23\pm 5^{+5}_{-4})\times 10^{-6}$ Belle-II:2023esi , which is about $2.7\sigma$ above the SM prediction $(4.16\pm 0.57)\times 10^{-6}$ Altmannshofer:2009ma ; Buras:2014fpa ; Becirevic:2023aov . Consequently, it has led to numerous studies exploring explanations beyond the SM, with possible excess sources may be sterile neutrinos Felkl:2023ayn ; Buras:2024ewl ; Datta:2023iln ; Rosauro-Alcaraz:2024mvx , axion-like particle Li:2024thq ; Li:2025ski ; Calibbi:2025rpx ; Altmannshofer:2024kxb , dark matter He:2022ljo ; He:2023bnk ; Altmannshofer:2023hkn ; Hou:2024vyw ; Fridell:2023ssf ; Gabrielli:2024wys ; Berezhnoy:2025osn ; MartinCamalich:2025srw ; Abdughani:2023dlr ; Wang:2023trd or modifications to the relevant Wilson coefficients from other NP Sumensari:2024sji ; Chen:2024jlj ; Allwicher:2023xba ; Athron:2023hmz ; Bause:2023mfe ; Chen:2024cll . The primary aim of this work is to derive constraints on a generic $Z^{\prime}$ portal dark matter model from rare meson decays with a neutrino pair in the final state and explore the complementarity between such decays and direct detection experiments for different DM–quark operator structures in the model.
@@
The structure of this paper is organized as follows. In section II, we establish the theoretical framework of a generic $Z^{\prime}$ portal model with all possible renormalizable couplings to quarks and DM. In section III, we present a systematic analysis of the rare meson decay processes, such as $B\to P(V)\chi\bar{\chi}$ and $K\to\pi\chi\bar{\chi}$. Using the latest experimental data, we derive joint constraints on the coupling strength and mass of light DM. In section IV, we compute the corresponding constraints on the cross section between light DM and nucleons, using the upper limits on the $Z^{\prime}$ portal model parameters derived from rare meson decays. By comparing these constraints with those from direct detection experiments, we assess the complementarity between these two approaches. Summary and conclusions are given in the last section V.
@@ II Theoretical Framework
where $f$ denotes all quarks in the SM. Note that $Z^{\prime}$ boson can also couple to leptons in the SM, corresponding to a leptophilic dark matter model. The collider experiments impose stringent constraints on the mass of the $Z^{\prime}$ boson if it couples simultaneously to both quarks and leptons Alves:2015pea . In this work, we mainly focus on the complementarity of rare meson decays to the parameter space of DM direct detection experiments. Therefore, we consider the scenario where dark matter interacts directly only with quarks via the $Z^{\prime}$ boson. This setup can be realized by assuming that the $U(1)$ charges of leptons are zero.
@@ III $B$ And $K$ Meson Decays
In this section, we derive bounds on the parameters of $Z^{\prime}$ portal model from the rare $B$ and $K$ meson decays. For convenience, we will use the dark low-energy effective field theory (DLEFT) Aebischer:2022wnl ; Song:2023jqm to describe these meson FCNC decays, and then match the simplified model described by Eq. (1) onto the DLEFT at the meson decay scales. Next, we will first briefly introduce the DLEFT and give the matching conditions. The analytic and numerical analysis of the differential branching ratio of $B$ and $K$ meson decays are also given in the end of this section.
@@ III.1 Effective Hamiltonian
As we know, the low-energy effective field theory(LEFT) Jenkins:2017jig is suitable for describing meson decays. The DLEFT is built on LEFT by adding new effective operators to describe the interactions between DM and SM sectors Aebischer:2022wnl ; Song:2023jqm . Since new physics effects arise from $Z^{\prime}$ in this work, it will only induce vector and axial-vector quark currents at low-energy scales. Restricting to dimension-6 operators, the effective Hamiltonian governing the $d_{i}\to d_{j}\chi\bar{\chi}$ transitions takes the form
@@
Here, $d$ denotes the down-type quarks under the mass basis, with $i$, $j$ the flavor indices. $\chi$ denotes the Fermion-type DM candidate.
@@
A renormalization scale $\mu$ is used to regularize the UV-divergnt loop contribution. We employ $\mu_{B}\approx 4.8$ GeV for $B$ meson decays and $\mu_{K}\approx 0.49$ GeV for $K$ meson decays.
@@
  $b\rightarrow s$ ${\cal B}(B^{+}\to K^{+}\nu\bar{\nu})$ $(4.16\pm 0.57)\times 10^{-6}$Altmannshofer:2009ma ; Buras:2014fpa ; Becirevic:2023aov $(23\pm 5^{+5}_{-4})\times 10^{-6}$Belle-II:2023esi
@@ III.2 $B$ Meson Decay
In this subsection, we will analyze the differential decay widths of rare $B$ meson decays, as well as their constraints on the parameter space of $Z^{\prime}$ portal model. The SM predictions and experimental measurements for the relevant meson decays are summarized in Table 1. Experimentally, neutrino pairs in the final states of these processes appear as missing energy. At quark level, transitions induced by the $b\to s$ or $b\to d$ decay with invisible final states could contribute to these processes, allowing these rare decays to constrain the couplings between light DM and quarks.
@@
Figure 2: Constraints on the effective couplings $g_{f}^{V(A)}g_{\chi}^{V(A)}$ versus DM mass $m_{\chi}$ from rare $B$ meson decays. The colored curves represent upper limits from different $b\to s(d)\chi\bar{\chi}$ transition channels, with the regions above each curve are experimentally excluded at $90\%$ confidence level.
Finally, taking into account the matching conditions given in Eq. (4), the parameters of $Z^{\prime}$ portal model can be directly constrained by the experimental measurements listed in Table 1, as illustrated in Figure 2. For the coupling product $g_{f}^{V}g_{\chi}^{V}$, the dominant constraint arises from the $B^{0}\to K^{0}\chi\bar{\chi}$, which is induced by the $b\to s$ transition.As shown in the figure, the constraints from the $B^{+}\to K^{+}\chi\bar{\chi}$ and $B^{0}\to K^{0}\chi\bar{\chi}$ channel are in good agreement. This consistency implies that, with higher expermental precision in the future, the neutral and charged processes will enable a valuable mutual cross-validation. As for the $b\to d$ transition, the most stringent constraint comes from the $B^{0}\to\pi^{0}\chi\bar{\chi}$ channel. In addition, as the DM mass $m_{\chi}$ approaches the kinematic endpoint $(m_{B}-m_{P,V})/2$, phase-space suppression rapidly degrades the constraints from a given decay channel. Hence, for $m_{\chi}>2.3~\mathrm{GeV}$, $B^{0}\to\pi^{0}\chi\bar{\chi}$ will provide a stronger constraint than $B^{0}\to K^{0}\chi\bar{\chi}$ channel due to the lighter final-state meson. For the coupling product $g_{f}^{V}g_{\chi}^{A}$, the constraints exhibit a similar trend as those for $g_{f}^{V}g_{\chi}^{V}$, with the main difference being an overall upward shift in the upper limits on the couplings. Since the axial-vector quark current yields no $B\to P$ hadron matrix elements, the constraints on $g_{f}^{A}g_{\chi}^{A}$ and $g_{f}^{A}g_{\chi}^{V}$ come exclusively from $B\to V$ modes, dominated by $B^{0}\to K^{*0}\chi\bar{\chi}$ decay channel.
@@ III.3 $K$ Meson Decay
In this subsection, we concentrate on the study of rare $K$ meson decays. Following the same approach as used for $B$-meson decays, we analyze their differential decay widths and derive corresponding constraints on the parameter space of the $Z^{\prime}$ portal model. Note that the Wilson coefficients induced by the $Z^{\prime}$ are almost pure real [^1] The imaginary part of the CKM factors is ignored in the present calculation., resulting in no contribution to the CP-violating $K_{L}\to\pi^{0}\chi\bar{\chi}$ process, whose amplitude is proportional to the imaginary part of the corresponding Wilson coefficients He:2022ljo . Therefore, we focuses on the $K^{+}\to\pi^{+}\chi\bar{\chi}$ decay channel. The hadronic matrix element for $K^{+}\to\pi^{+}$ transitions is dominated by the vector quark current. In the present work, we adopt the parameterization of this hadronic matrix element as established in Refs. Kamenik:2011vy ; He:2022ljo , given by
@@
Herein, we elaborate on the constraints on the parameter space of the $Z^{\prime}$ portal model imposed by $K$ meson decays, As shown in Figure 3, the exclusion limits for the coupling product $g_{f}^{V}g_{\chi}^{V(A)}$ are found to be on the order of $10^{-1}$. Notably, the constraint associated with $g_{f}^{V}g_{\chi}^{A}$ is marginally more stringent, and this disparity stems from the distinct decay widths of the relevant $K$ meson channels.
@@ IV Complementarity of rare meson decays with DM direct detection
DM direct detection experiments have been primarily dedicated to searching for WIMPs, with mass between GeV to TeV scale. However, their sensitivity to light DM at the GeV scale and below is significantly weaker. This is due to the kinetic energy of light DM particles is small and very little gets transferred to a nucleon in elastic scattering. In this section, we use the upper bounds on the $Z^{\prime}$ portal model parameters derived from rare meson decays in the previous section to constrain the scattering cross section of light dark matter off nucleons.
@@
DM direct detection experiments are broadly categorized into spin-independent (SI) and spin-dependent (SD) searches. For SI interactions, the scattering cross section scales with the square of the target nucleon number $A$, which motivates the use of heavy nuclei to enhance the signal. Therefore, SI searches generally place stronger constraints on the cross section than SD searches. In the $Z^{\prime}$ portal model, the induced operators involving a vector quark current, i.e. $\mathcal{O}_{d\chi}^{VV}$ and $\mathcal{O}_{d\chi}^{VA}$, correspond to spin-independent interactions, while the operators with an axial-vector quark current, i.e. $\mathcal{O}_{d\chi}^{AV}$ and $\mathcal{O}_{d\chi}^{AA}$, give rise to spin-dependent interactions. The DM-nucleon cross section induced by operators $\mathcal{O}_{d\chi}^{VA}$ and $\mathcal{O}_{d\chi}^{AV}$ depend on the momentum exchange An:2012va . Given the typical velocity of the DM in the galactic halo $v\sim 10^{-3}$, these cross sections are highly suppressed by a factor of $v^{2}$. The properties of these operators are collected in Table 2. Next, we focus on analyzing the constraints on the DM-nucleon cross section associated with operators $\mathcal{O}_{d\chi}^{VV}$ and $\mathcal{O}_{d\chi}^{AA}$. Using the joint constraints on the DM mass and coupling products from rare meson decays, both SI and SD DM-nucleon cross section can be calculated by package micrOMEGAs Belanger:2020gnr .
@@
Figure 5 shows the constraints on the spin-dependent DM-neutron(proton) cross section from rare $B$ and $K$ meson decays. For comparison, constraints from the LUX, XENON1T, XENONnT, PICO-60, and LZ experiments on the spin-dependent DM-neutron(proton) cross section are also shown in the figure. Among these experiments, the LZ experiment provides the strongest constraints for DM masses above 10 GeV for both spin-dependent DM-neutron and DM-proton cross sections. For DM mass in the few-GeV region, the PICO-60 experiment places the strongest constraints for spin-dependent DM-proton cross section. The solid lines in the Figure 5 provide constraints on the spin-dependent DM neutron(proton) cross section derived from the rare meson decays. Among these, the most stringent limit is placed by $B^{0}\to K^{*0}\chi\bar{\chi}$, which is on the order of $10^{-42}\,\text{cm}^{2}$ to $10^{-40}\,\text{cm}^{2}$ for light DM in the mass range 100 MeV to 1 GeV. The solid lines in Figure 5 can also roughly be extended to represent the SD momentum-dependent interaction scenario.

### v2

@@ Abstract
Current dark matter direct detection experiments have low sensitivity to sub-GeV dark matter. In this work, we demonstrate that rare $B$ and $K$ meson decays with missing energy in the final state can serve as efficient probes in this mass range. We analyze a generic $Z^{\prime}$ portal dark matter model and derive upper limits on its parameters from experimental bounds on the rare $B$ and $K$ meson decays. Our results show that such meson decay processes provide complementary constraints to current direct detection experiments for sub-GeV dark matter, particularly for interaction forms mediated by dark matter momentum-dependent operators.
@@ I Introduction
Abundant astronomical and cosmological evidence indicates that about a quarter of the matter in the Universe is composed of dark matter (DM) ParticleDataGroup:2024cfk. However, the nature of DM remains unknown as the Standard Model (SM) of particle physics fails to provide a viable DM candidate. The most widely studied DM candidates are the so-called weakly interacting massive particles (WIMPs), which could naturally provide the right amount of DM relic density, a coincidence known as the WIMP miracle. Most current DM direct detection experiments target the search for WIMPs, with mass range roughly on the order of GeV to TeV. DM direct detection experiments mostly aim to observe elastic or inelastic scattering of DM particles with atomic nuclei, such as liquid xenon detectors XENONnT XENON:2023cxc, LUX-ZEPLIN LZ:2022lsv, and PandaX-4T PandaX:2022xas. While no confirmed DM signal has been observed, these experiments have placed stringent limits on the DM–nucleon scattering cross section. However, their sensitivity drops sharply for DM masses below the GeV scale, as the nuclear recoil signals from its elastic scattering with nuclei typically lie below the detection energy threshold and are overwhelmed by background noise. Searching for light DM in heavy meson decays can serve as a complement to DM direct detection, and its sensitivity in some models may be higher than that of direct detection experiments.
@@
Semileptonic flavor-changing neutral current (FCNC) decays are highly suppressed in the SM due to their loop origin and the Glashow-Iliopoulos–Maiani (GIM) mechanism, making them exceptionally sensitive to new physics (NP) effects. In particular, decays with a neutrino pair in the final state provide a distinct probe of NP compared to those with charged leptons. For semileptonic decays to charged leptons, NP effects are primarily probed through modifications of the relevant Wilson coefficients, mediated by virtual NP particles. In contrast, for decays to a neutrino pair, NP particles themselves can exist as missing energy hiding in the final state, since the neutrinos are not detected in experiments. In addition to the SM neutrinos, any electrically neutral, undetectable particle can manifest as missing energy in such processes. Notably, the Belle II experiment has recently reported the first evidence of $B^{+}\to K^{+}\nu\bar{\nu}$ with the branching ratio of $\text{Br}(B^{+}\to K^{+}\nu\bar{\nu})=(23\pm 5^{+5}_{-4})\times 10^{-6}$ Belle-II:2023esi, which is about $2.7\sigma$ above the SM prediction $(4.16\pm 0.57)\times 10^{-6}$ Hou:2024vyw; Altmannshofer:2009ma; Buras:2014fpa; Becirevic:2023aov. Consequently, it has led to numerous studies exploring explanations beyond the SM, with possible excess sources including sterile neutrinos Felkl:2023ayn; Buras:2024ewl; Datta:2023iln; Rosauro-Alcaraz:2024mvx, axion-like particle Li:2024thq; Li:2025ski; Calibbi:2025rpx; Altmannshofer:2024kxb, dark matter He:2022ljo; He:2023bnk; He:2024iju; Aliev:2025hyp; He:2025jfc; Altmannshofer:2023hkn; Hou:2024vyw; Fridell:2023ssf; Gabrielli:2024wys; Berezhnoy:2025osn; MartinCamalich:2025srw; Abdughani:2023dlr; Wang:2023trd or modifications to the relevant Wilson coefficients from other NP Sumensari:2024sji; Chen:2024jlj; Allwicher:2023xba; Athron:2023hmz; Bause:2023mfe; Chen:2024cll. The primary aim of this work is to derive constraints on a generic $Z^{\prime}$ portal dark matter model from rare meson decays with a neutrino pair in the final state and explore the complementarity between such decays and direct detection experiments for different DM–quark operator structures in the model.
@@
This paper is organized as follows. In section II, we establish the theoretical framework of a generic $Z^{\prime}$ portal model with all possible renormalizable couplings to quarks and DM. In section III, we present a systematic analysis of the rare meson decay processes, such as $B\to P(V)\chi\bar{\chi}$ and $K\to\pi\chi\bar{\chi}$. Using the latest experimental data, we derive joint constraints on the coupling strength and mass of light DM. In section IV, we compute the corresponding constraints on the cross section between light DM and nucleons, using the upper limits on the $Z^{\prime}$ portal model parameters derived from rare meson decays. By comparing these constraints with those from direct detection experiments, we assess the complementarity between these two approaches. Summary and conclusions are given in the last section V.
@@ II Theoretical Framework
where $f$ denotes all quarks in the SM. Note that the $Z^{\prime}$ boson can also couple to leptons in the SM, corresponding to a leptophilic dark matter model. The collider experiments impose stringent constraints on the mass of the $Z^{\prime}$ boson if it couples simultaneously to both quarks and leptons Alves:2015pea. In this work, we mainly focus on the complementarity of rare meson decays to the parameter space of DM direct detection experiments. Therefore, we consider the scenario where dark matter interacts directly only with quarks via the $Z^{\prime}$ boson. This setup can be realized by assuming that the $U(1)$ charges of leptons are zero.
@@ III $B$ And $K$ Meson Decays
In this section, we derive bounds on the parameters of the $Z^{\prime}$ portal model from the rare $B$ and $K$ meson decays. For convenience, we will use the dark low-energy effective field theory (DLEFT) Aebischer:2022wnl; Song:2023jqm to describe these meson FCNC decays, and then match the simplified model described by Eq. (1) onto the DLEFT at the meson decay scales. Next, we will first briefly introduce the DLEFT and give the matching conditions. The analytic and numerical analysis of the differential branching ratio of $B$ and $K$ meson decays is given in the end of this section.
@@ III.1 Effective Hamiltonian
As we know, the low-energy effective field theory(LEFT) Jenkins:2017jig; Liang:2023yta is suitable for describing meson decays. The DLEFT is built on LEFT by adding new effective operators to describe the interactions between DM and SM sectors Aebischer:2022wnl; Song:2023jqm. Since new physics effects arise from $Z^{\prime}$ in this work, it will only induce vector and axial-vector quark currents at low-energy scales. Restricting to dimension-6 operators, the effective Hamiltonian governing the $d_{i}\to d_{j}\chi\bar{\chi}$ transitions takes the form
@@
Here, $d$ denotes the down-type quarks in the mass basis, with $i$, $j$ the flavor indices. $\chi$ denotes the fermion-type DM candidate.
@@
A renormalization scale $\mu$ is used to regularize the UV-divergent loop contribution. We employ $\mu_{B}\approx 4.8$ GeV for $B$ meson decays and $\mu_{K}\approx 0.49$ GeV for $K$ meson decays.
@@
  $b\rightarrow s$ ${\cal B}(B^{+}\to K^{+}\nu\bar{\nu})$ $(4.16\pm 0.57)\times 10^{-6}$Hou:2024vyw; Altmannshofer:2009ma; Buras:2014fpa; Becirevic:2023aov $(23\pm 5^{+5}_{-4})\times 10^{-6}$Belle-II:2023esi
@@ III.2 $B$ Meson Decay
In this subsection, we will analyze the differential decay widths of rare $B$ meson decays, as well as their constraints on the parameter space of the $Z^{\prime}$ portal model. The SM predictions and experimental measurements for the relevant meson decays are summarized in Table 1. Experimentally, neutrino pairs in the final states of these processes appear as missing energy. At quark level, transitions induced by the $b\to s$ or $b\to d$ decay with invisible final states could contribute to these processes, allowing these rare decays to constrain the couplings between light DM and quarks.
@@
Figure 2: Constraints on the effective couplings $g_{f}^{V(A)}g_{\chi}^{V(A)}$ versus DM mass $m_{\chi}$ from rare $B$ meson decays. The colored curves represent upper limits from different $b\to s(d)\chi\bar{\chi}$ transition channels, with the regions above each curve experimentally excluded at $90\%$ confidence level.
Finally, taking into account the matching conditions given in Eq. (4), the parameters of the $Z^{\prime}$ portal model can be directly constrained by the experimental measurements listed in Table 1, as illustrated in Figure 2. For the coupling product $g_{f}^{V}g_{\chi}^{V}$, the dominant constraint arises from the $B^{0}\to K^{0}\chi\bar{\chi}$, which is induced by the $b\to s$ transition.As shown in the figure, the constraints from the $B^{+}\to K^{+}\chi\bar{\chi}$ and $B^{0}\to K^{0}\chi\bar{\chi}$ channel are in good agreement. This consistency implies that, with higher experimental precision in the future, the neutral and charged processes will enable a valuable mutual cross-validation. As for the $b\to d$ transition, the most stringent constraint comes from the $B^{0}\to\pi^{0}\chi\bar{\chi}$ channel. In addition, as the DM mass $m_{\chi}$ approaches the kinematic endpoint $(m_{B}-m_{P,V})/2$, phase-space suppression rapidly degrades the constraints from a given decay channel. Hence, for $m_{\chi}>2.3~\mathrm{GeV}$, $B^{0}\to\pi^{0}\chi\bar{\chi}$ will provide a stronger constraint than $B^{0}\to K^{0}\chi\bar{\chi}$ channel due to the lighter final-state meson. For the coupling product $g_{f}^{V}g_{\chi}^{A}$, the constraints exhibit a similar trend as those for $g_{f}^{V}g_{\chi}^{V}$, with the main difference being an overall upward shift in the upper limits on the couplings. Since the axial-vector quark current yields no $B\to P$ hadron matrix elements, the constraints on $g_{f}^{A}g_{\chi}^{A}$ and $g_{f}^{A}g_{\chi}^{V}$ come exclusively from $B\to V$ modes, dominated by $B^{0}\to K^{*0}\chi\bar{\chi}$ decay channel.
@@ III.3 $K$ Meson Decay
In this subsection, we concentrate on the study of rare $K$ meson decays. Following the same approach as used for $B$-meson decays, we analyze their differential decay widths and derive corresponding constraints on the parameter space of the $Z^{\prime}$ portal model. Note that the Wilson coefficients induced by the $Z^{\prime}$ are almost purely real [^1] The imaginary part of the CKM factors is ignored in the present calculation., resulting in no contribution to the CP-violating $K_{L}\to\pi^{0}\chi\bar{\chi}$ process, whose amplitude is proportional to the imaginary part of the corresponding Wilson coefficients He:2022ljo. Therefore, we focus on the $K^{+}\to\pi^{+}\chi\bar{\chi}$ decay channel. The hadronic matrix element for $K^{+}\to\pi^{+}$ transitions is dominated by the vector quark current. In the present work, we adopt the parameterization of this hadronic matrix element as established in Refs. Kamenik:2011vy; He:2022ljo, given by
@@
Herein, we elaborate on the constraints on the parameter space of the $Z^{\prime}$ portal model imposed by $K$ meson decays. As shown in Figure 3, the exclusion limits for the coupling product $g_{f}^{V}g_{\chi}^{V(A)}$ are found to be on the order of $10^{-1}$. Notably, the constraint associated with $g_{f}^{V}g_{\chi}^{A}$ is marginally more stringent, and this disparity stems from the distinct decay widths of the relevant $K$ meson channels.
@@ IV Complementarity of rare meson decays with DM direct detection
DM direct detection experiments have been primarily dedicated to searching for WIMPs, with masses from GeV to TeV scale. However, their sensitivity to light DM at the GeV scale and below is significantly weaker. This is because the kinetic energy of light DM particles is small and very little gets transferred to a nucleon in elastic scattering. In this section, we use the upper bounds on the $Z^{\prime}$ portal model parameters derived from rare meson decays in the previous section to constrain the scattering cross section of light dark matter off nucleons.
@@
DM direct detection experiments are broadly categorized into spin-independent (SI) and spin-dependent (SD) searches. For SI interactions, the scattering cross section scales with the square of the target nucleon number $A$, which motivates the use of heavy nuclei to enhance the signal. Therefore, SI searches generally place stronger constraints on the cross section than SD searches. In the $Z^{\prime}$ portal model, the induced operators involving a vector quark current, i.e. $\mathcal{O}_{d\chi}^{VV}$ and $\mathcal{O}_{d\chi}^{VA}$, correspond to spin-independent interactions, while the operators with an axial-vector quark current, i.e. $\mathcal{O}_{d\chi}^{AV}$ and $\mathcal{O}_{d\chi}^{AA}$, give rise to spin-dependent interactions. The DM-nucleon cross section induced by operators $\mathcal{O}_{d\chi}^{VA}$ and $\mathcal{O}_{d\chi}^{AV}$ depend on the momentum exchange An:2012va. Given the typical velocity of the DM in the galactic halo $v\sim 10^{-3}$, these cross sections are highly suppressed by a factor of $v^{2}$. The properties of these operators are collected in Table 2. Next, we focus on analyzing the constraints on the DM-nucleon cross section associated with operators $\mathcal{O}_{d\chi}^{VV}$ and $\mathcal{O}_{d\chi}^{AA}$. Using the joint constraints on the DM mass and coupling products from rare meson decays, both SI and SD DM-nucleon cross sections can be calculated by package micrOMEGAs Belanger:2020gnr.
@@
Figure 5 shows the constraints on the spin-dependent DM-neutron(proton) cross section from rare $B$ and $K$ meson decays. For comparison, constraints from the LUX, XENON1T, XENONnT, PICO-60, and LZ experiments on the spin-dependent DM-neutron(proton) cross section are also shown in the figure. Among these experiments, the LZ experiment provides the strongest constraints for DM masses above 10 GeV for both spin-dependent DM-neutron and DM-proton cross sections. For DM mass in the few-GeV region, the PICO-60 experiment places the strongest constraints for spin-dependent DM-proton cross section. The solid lines in Figure 5 provide constraints on the spin-dependent DM-neutron(proton) cross section derived from the rare meson decays. Among these, the most stringent limit is placed by $B^{0}\to K^{*0}\chi\bar{\chi}$, which is on the order of $10^{-42}\,\text{cm}^{2}$ to $10^{-40}\,\text{cm}^{2}$ for light DM in the mass range 100 MeV to 1 GeV. The solid lines in Figure 5 can also roughly be extended to represent the SD momentum-dependent interaction scenario.